# Discrete Diffusion-Based Model-Level Explanation of Heterogeneous GNNs with Node Features

### v1

@@ 2. Node feature generation
We conduct experiments on seven datasets for node and graph classification (Fu et al. 2020; Morris et al. 2020; Ying et al. 2019; Wang et al. 2021)—three real-world datasets (DBLP, IMDB, MUTAG) and four synthetic datasets (BA-Shapes, Tree-Cycle, Tree-Grids, BA-3Motif). DBLP and IMDB have actual node features, the other datasets not.
@@
We use a high-performing GNN architecture for each dataset: GraphSAGE (Hamilton, Ying, and Leskovec 2017) for DBLP, HAN (Wang et al. 2019) for IMDB, and three-layer GCNs for MUTAG and the synthetic datasets. To prevent overfitting, we apply early stopping with a patience of 100 epochs, halting training if the validation score does not improve within this window.
@@
We compare DiGNNExplainer to the following state-of-the-art model-level explanation methods: (1) XGNN (Yuan et al. 2020), which leverages reinforcement learning to generate explanations. We extend the original implementation to support node classification tasks. (2) GNNInterpreter (Wang and Shen 2023), which improves upon XGNN by being computationally efficient. We extend it for node classification by creating class-specific embeddings using average node pooling. (3) D4Explainer (Chen et al. 2023), which employs a diffusion-based approach to generate model-level explanations for both node and graph classification tasks.
@@
We compare our discrete diffusion-based graph generator against two baselines: Our first baseline is a variational autoencoder-based model that jointly learns the graph structure and node features. Our second baseline replaces our discrete diffusion-based node feature generator DiTabDDPM with the continuous diffusion-based generator TabDDPM. Regarding the former, for MUTAG and the synthetic datasets, we use the original implementation of GraphVAE [^4] https://github.com/deepfindr/gvae (Simonovsky and Komodakis 2018) using only node types as node features. For DBLP and IMDB, we extend GraphVAE to generate explanation graphs with actual node features by concatenating one-hot encoded node types to actual node features. We omit GraphVAE’s support for node order invariance as it does not scale to graphs of our size. As GraphVAE requires all nodes to have the same number of features, but DBLP has a varying number of features per node type, we select a feature size of 50 (using feature selection) for all node types of DBLP, to be able to compare all our DBLP experiments with the baseline. For both the VAE and TabDDPM baselines, which always operate in a continuous feature space, we discretize features that are supposed to be discrete (e.g., bag-of-word features) via a threshold.
@@
To assess faithfulness, following (Agarwal et al. 2022), we assess both predictive faithfulness $\mathit{PF}$ and ground-truth faithfulness $\mathit{GF}$: Regarding the former, we compute the average class score across all classes as $\mathit{PF}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{C}\sum_{i=1}^{C}P({G_{i}})$ where $C$ is the number of classes, $\mathcal{G}_{i}$ is our explanation graph for class $i$, and $P(G_{i})$ is the GNN’s predicted probability for class $i$ on explanation graph $G_{i}$. A higher score indicates greater alignment between the explanation graphs and the GNN’s original decision. Regarding the latter, we compute ground-truth faithfulness as follows: For each class, we assume that we have a set of ground-truth explanation motifs (which are readily available for synthetic datasets and can be obtained via community detection for real-world datasets as described in the appendix). Then we define $\mathit{GF}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{C}\sum_{i=1}^{C}\frac{1}{M_{i}}\sum_{j=1}^{M_{i}}\mathds{1}_{(M_{i,j}\subseteq G_{i})}$ where $M_{i}$ is the number of available ground-truth motifs for class $i$, $G_{i}$ is the explanation graph for class $i$, and $\mathds{1}_{(M_{i,j}\subseteq G_{i})}$ is the indicator function that equals 1 if the motif $M_{i,j}$, is a subgraph of the explanation graph, and 0 otherwise. A higher ground-truth faithfulness score indicates that our generated explanations are closer to the ground-truth explanations.
@@
For the node classification datasets DBLP, IMDB, BA-Shapes, Tree-Cycle, Tree-Grids, we synthesize 1,536 graphs per dataset, some of which may be discarded as they are inconsistent to the metagraph. We use 200 graphs to train the GNN, splitting them into 128 for training, 32 for validation, and 40 for testing the graph generation model. For MUTAG (58 graphs), we allocate 38 for training, 9 for validation, and 11 for testing. For BA-3Motif (323 graphs), the split is 206 for training, 52 for validation, and 65 for testing. To generate candidate explanation graphs, we use our graph generation model to produce 256 graphs for each of the 6 graph sizes (either 10 to 15 nodes or 5 to 10 nodes), resulting in 1,536 candidate graphs in total.
Table 1: Comparisons for average MMD distances between explanation graphs of each class and the real-world graphs, for Node Type Distributions (NTD), Node Degree distributions (Deg.), clustering coefficients (Clus.), spectrum distributions (Spec.).
@@
Table 2: Table showing evaluation results for feature selection experiments using variance for DBLP and IMDB. This includes average Predictive Faithfulness (PF) and Ground-truth Faithfulness (GF) for 10 runs, average cosine similarity for node features (Node feat. (cos)). We include a comparison between our new DiTabDDPM and the original TabDDPM.
@@
DiTabDDPM (50 feat.) 0.999 0.882 0.229
DiTabDDPM (8 discr. feat.) 0.835 0.794 0.353
DiTabDDPM (4 discr. feat.) 0.818 0.809 0.386
DiTabDDPM (2 discr. feat.) 0.975 0.817 0.397
TabDDPM (50 feat.) 0.993 0.829 0.196
TabDDPM (8 discr. feat.) 0.995 0.822 0.221
TabDDPM (4 discr. feat.) 0.965 0.825 0.232
TabDDPM (2 discr. feat.) 0.979 0.827 0.235
VAE (baseline) 0.655 0.717 0.0
@@
DiTabDDPM (3066 feat., all) 0.955 0.890 0.029
DiTabDDPM (10 discr. feat.) 0.924 0.900 0.083
DiTabDDPM (5 discr. feat.) 0.954 0.883 0.069
DiTabDDPM (2 discr. feat.) 0.976 0.910 0.040
TabDDPM (3066 feat., all) 0.912 0.890 0.048
TabDDPM (10 discr. feat.) 0.954 0.903 0.016
TabDDPM (5 discr. feat.) 0.988 0.909 0.011
TabDDPM (2 discr. feat.) 0.974 0.913 0.005
VAE (baseline) 0.999 0.360 0.0
Realism of Graph Structure
Implementation (1) We generate graphs that serve as explanations. In our experiments, we restrict the explanation graph size to a range of 5 to 15 nodes to ensure human-intelligible explanation graphs. (2) We compare the average MMD (Table 1) between the explanation graphs of each class and the real graphs for node type distribution, node degree, clustering coefficient, spectrum distributions for 50 graphs.
Results DiGNNExplainer achieves the highest similarity for node degree distribution across all datasets except BA-Shapes, and the highest similarity for node type distribution (NTD) and clustering coefficient (Clus.) for the real-world datasets (Table 1).
Interpretation DiGNNExplainer generates the most realistic graphs, preserving the distributions of graphs for all datasets and node types for heterogeneous datasets. The homogeneous datasets have a single node type, hence the comparison always yields a score of 0.0.
Realism of Node Features
In this section, we identify the approach for generating the most realistic node features using cosine similarity (Node feat.(cos) in Table 2). We compare DiTabDDPM and TabDDPM for node feature realism for various node feature sizes. The notations used for our feature selection experiments in the Table 2 are defined as: (i) DiTabDDPM - Discrete (bag-of-words) node features generated using DiTabDDPM and continuous node features generated using TabDDPM (ii) TabDDPM - All node features generated using TabDDPM (iii) DiTabDDPM (8 discr. feat.) indicates that the node to be classified has a feature size of 8.
Implementation (1) For discrete node features, we use one-hot encoding of word frequencies, i.e., in the range of 0 to 5 for Paper nodes. (2) For VAE, we learn the graphs along with node features which include all features, i.e., 50 for DBLP and 3066 for IMDB. In VAE, we average the cosine similarity of the node features of the node to be classified for 50 graphs.
Results Cosine similarity (Node feat.) results are higher for DiTabDDPM compared to TabDDPM, with the highest score of 0.397 for DBLP.
Interpretation DiTabDDPM, which learns a discrete space for discrete features, generates more realistic node features compared to the original TabDDPM, which learns a continuous space for discrete features.
Faithfulness of Explanations
Table 3: Comparisons for Predictive Faithfulness (PF) and Ground-truth Faithfulness (GF) with existing approaches.
@@
DiGNNExplainer 0.895 0.882 0.947 0.913 1.0 0.425 0.315 0.766 0.525 0.716 0.507 0.893 0.432 0.639
@@
We compare the faithfulness of our approach for - (1) Existing model-level explanation approaches (Table 3) that do not utilize actual node features. For this, we also run experiments on the datasets used in the original baseline explainer implementation. (2) Heterogeneous datasets that use node features of varying sizes generated using TabDDPM and DiTabDDPM (Table 2), to observe the impact of node feature size on GNN predictions (similarly done in existing work (Dong et al. 2021)), and explanation quality. We use the same generated graphs and the same ground truth motifs for our approaches with different node feature sizes.
@@
Implementation (1) The baseline implementations do not support node features and heterogeneous data. Hence, to implement baselines on our heterogeneous datasets DBLP and IMDB, we are forced to remove node features and use one-hot encoding of node types. Our baseline explainers support homogeneous datasets and are implemented using a GCN model. We use the original model architecture for each of the baselines. We convert the datasets to homogeneous and train a GCN model on the homogeneous datasets. (2) Same as in DiGNNExplainer, we generate graphs in our baselines for each node size using separate models. The baseline approaches allow setting a maximum node size for graph generation. (3) We use early stopping for all trained GNNs to prevent overfitting.
Results (1) In Table 3, DiGNNExplainer, which uses node features, outperforms all baseline explainers on ground-truth faithfulness score for all datasets and generates the highest predictive faithfulness scores for all real-world datasets. (2) In Table 2, for DBLP, we obtain the best explanation graphs with a maximum ground-truth faithfulness score of 0.882 for DBLP using DiTabDDPM (50 feat.) with 50 features. For IMDB, we obtain similar maximum scores of 0.91 for DiTabDDPM (2 discr. feat.) and TabDDPM (2 discr. feat.) with 2 features. We determine the best explanation using the ground-truth faithfulness score rather than only the predictive faithfulness score because models with a high quality of explanation can be trusted.
@@
Interpretation (1) The explanations generated for our real-world datasets (Table 3) using DiGNNExplainer are faithful, as we obtained both high predictive faithfulness and ground-truth faithfulness scores. (2) We obtained a lower ground-truth faithfulness score for our baseline explainers due to their inability to model heterogeneous structures and lack of node features. (3) The reason for a low predictive faithfulness score for synthetic datasets is that the 3-layer GCN model predicts the node class of base Barabási–Albert(BA) graph accurately compared to all other classes (house, cycle, grid). (4) The reason for the low ground-truth faithfulness score for VAE is that graph generation via VAE is unable to preserve the underlying heterogeneous structures. (5) The reason for the high predictive faithfulness score in VAE for IMDB is that the GNN is trained with all features of the dataset, where all node types except Movie have continuous features, and the continuous features generated using VAE are more realistic. (6) Realistic discrete features generated by learning a discrete space in DiTabDDPM result in highly faithful explanations, but we also observe that the node features for faithful explanations are not always the most realistic, i.e., do not have the highest cosine similarity values. (7) The DBLP experiment (Table 2) with a large feature size produces both high predictive faithfulness and ground-truth faithfulness scores, indicating that the GNN model can be trusted.
From the result interpretations (5, 6, 7) above, we can infer that node feature generation technique influences the quality of explanation graphs.
Experiment on Feature Selection

### v2

@@ 4.1. Evaluation Setup
We conduct experiments on seven datasets for node and graph classification (Fu et al., 2020; Morris et al., 2020; Ying et al., 2019; Wang et al., 2021)—three real-world datasets (DBLP, IMDB, MUTAG) and four synthetic datasets (BA-Shapes, Tree-Cycle, Tree-Grids, BA-3Motif) in Table 7. DBLP and IMDB have actual node features, the other datasets not.
@@
We use a high-performing GNN architecture for each dataset: GraphSAGE (Hamilton et al., 2017) for DBLP, HAN (Wang et al., 2019) for IMDB, and three-layer GCNs for MUTAG and the synthetic datasets. To prevent overfitting, we apply early stopping with a patience of 100 epochs, halting training if the validation score does not improve within this window.
@@
We compare DiGNNExplainer to the following state-of-the-art model-level explanation methods: (1) XGNN (Yuan et al., 2020), which leverages reinforcement learning to generate explanations. We extend the original implementation to support node classification tasks. (2) GNNInterpreter (Wang and Shen, 2023), which improves upon XGNN by being computationally efficient. We extend it for node classification by creating class-specific embeddings using average node pooling. (3) D4Explainer (Chen et al., 2023), which employs a diffusion-based approach to generate model-level explanations for both node and graph classification tasks.
@@
We compare our discrete diffusion-based graph generator against two baselines: Our first baseline is a variational autoencoder-based model that jointly learns the graph structure and node features. Our second baseline replaces our discrete diffusion-based node feature generator DiTabDDPM with the continuous diffusion-based generator TabDDPM. Regarding the former, for MUTAG and the synthetic datasets, we use the original implementation of GraphVAE [^4] https://github.com/deepfindr/gvae (Simonovsky and Komodakis, 2018) using only node types as node features. For DBLP and IMDB, we extend GraphVAE to generate explanation graphs with actual node features by concatenating one-hot encoded node types to actual node features. We omit GraphVAE’s support for node order invariance as it does not scale to graphs of our size. As GraphVAE requires all nodes to have the same number of features, but DBLP has a varying number of features per node type, we select a feature size of 50 (using feature selection) for all node types of DBLP, to be able to compare all our DBLP experiments with the baseline. For both the VAE and TabDDPM baselines, which always operate in a continuous feature space, we discretize features that are supposed to be discrete (e.g., bag-of-word features) via a threshold.
@@
To assess faithfulness, following (Agarwal et al., 2022), we assess both predictive faithfulness $\mathit{PF}$ and ground-truth faithfulness $\mathit{GF}$: Regarding the former, we compute the average class score across all classes as $\mathit{PF}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{C}\sum_{i=1}^{C}P({G_{i}})$ where $C$ is the number of classes, $\mathcal{G}_{i}$ is our explanation graph for class $i$, and $P(G_{i})$ is the GNN’s predicted probability for class $i$ on explanation graph $G_{i}$. A higher score indicates greater alignment between the explanation graphs and the GNN’s original decision. Regarding the latter, we compute ground-truth faithfulness as follows: For each class, we assume that we have a set of ground-truth explanation motifs (which are readily available for synthetic datasets and can be obtained via community detection for real-world datasets as described in the appendix). Then we define $\mathit{GF}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{C}\sum_{i=1}^{C}\frac{1}{M_{i}}\sum_{j=1}^{M_{i}}\mathds{1}_{(M_{i,j}\subseteq G_{i})}$ where $M_{i}$ is the number of available ground-truth motifs for class $i$, $G_{i}$ is the explanation graph for class $i$, and $\mathds{1}_{(M_{i,j}\subseteq G_{i})}$ is the indicator function that equals 1 if the motif $M_{i,j}$, is a subgraph of the explanation graph, and 0 otherwise. A higher ground-truth faithfulness score indicates that our generated explanations are closer to the ground-truth explanations.
@@
For the node classification datasets DBLP, IMDB, BA-Shapes, Tree-Cycle, and Tree-Grids, we synthesize 1,536 graphs per dataset, some of which may be discarded as they are inconsistent with the metagraph. We use 200 graphs to train the GNN, splitting them into 128 for training, 32 for validation, and 40 for testing the graph generation model. For MUTAG (58 graphs), we allocate 38 for training, 9 for validation, and 11 for testing. For BA-3Motif (323 graphs), the split is 206 for training, 52 for validation, and 65 for testing. To generate candidate explanation graphs, we use our graph generation model to produce 256 graphs for each of the 6 graph sizes (either 10 to 15 nodes or 5 to 10 nodes), resulting in 1,536 candidate graphs in total.
Table 1. Comparisons for average MMD distances between explanation graphs of each class and the reference graphs, for node type distributions (NTD), node degree distributions (Deg.), clustering coefficients (Clus.), spectrum distributions (Spec.).
@@
Table 2. Feature selection experiments for DBLP and IMDB. We report the average cosine similarity for node features (Node feat. (cos)), and the mean and variance of predictive faithfulness (PF) and ground-truth faithfulness (GF) over 10 runs. We compare our proposed DiTabDDPM with the original TabDDPM and the VAE baseline.
@@
DiTabDDPM (50 feat.) 0.229 0.999 $\pm$ 0.0 0.882 $\pm$ 0.029
DiTabDDPM (8 discr. feat.) 0.353 0.835 $\pm$ 0.090 0.794 $\pm$ 0.052
DiTabDDPM (4 discr. feat.) 0.386 0.818 $\pm$ 0.056 0.809 $\pm$ 0.051
DiTabDDPM (2 discr. feat.) 0.397 0.975 $\pm$ 0.036 0.817 $\pm$ 0.048
TabDDPM (50 feat.) 0.196 0.993 $\pm$ 0.005 0.829 $\pm$ 0.058
TabDDPM (8 discr. feat.) 0.221 0.995 $\pm$ 0.002 0.822 $\pm$ 0.052
TabDDPM (4 discr. feat.) 0.232 0.965 $\pm$ 0.046 0.825 $\pm$ 0.086
TabDDPM (2 discr. feat.) 0.235 0.979 $\pm$ 0.037 0.827 $\pm$ 0.032
VAE (baseline) 0.0 0.655 $\pm$ 0.042 0.717 $\pm$ 0.046
@@
DiTabDDPM (3066 feat., all) 0.029 0.955 $\pm$ 0.084 0.890 $\pm$ 0.029
DiTabDDPM (10 discr. feat.) 0.083 0.924 $\pm$ 0.091 0.900 $\pm$ 0.042
DiTabDDPM (5 discr. feat.) 0.069 0.954 $\pm$ 0.073 0.883 $\pm$ 0.030
DiTabDDPM (2 discr. feat.) 0.040 0.976 $\pm$ 0.065 0.910 $\pm$ 0.051
TabDDPM (3066 feat., all) 0.048 0.912 $\pm$ 0.104 0.890 $\pm$ 0.036
TabDDPM (10 discr. feat.) 0.016 0.954 $\pm$ 0.073 0.903 $\pm$ 0.034
TabDDPM (5 discr. feat.) 0.011 0.988 $\pm$ 0.019 0.909 $\pm$ 0.042
TabDDPM (2 discr. feat.) 0.005 0.974 $\pm$ 0.065 0.913 $\pm$ 0.039
VAE (baseline) 0.0 0.999 $\pm$ 0.0 0.360 $\pm$ 0.165
Table 3. Comparison of explainers in terms of predictive faithfulness (PF) and ground-truth faithfulness (GF).
@@
DiGNNExplainer-GCN 0.641 0.800 0.410 0.850 1.0 0.425 0.315 0.766 0.525 0.716 0.507 0.893 0.432 0.639
DiGNNExplainer 0.999 0.882 0.974 0.913 1.0 0.425 0.315 0.766 0.525 0.716 0.507 0.893 0.432 0.639
4.2. Realism of Graph Structure
Implementation
(1) We generate graphs that serve as explanations. In our experiments, we restrict the explanation graph size to a range of 5 to 15 nodes to ensure human-intelligible explanation graphs. (2) We compare the average MMD (Table 1) between the explanation graphs of each class and the real graphs for node type distribution, node degree, clustering coefficient, and spectrum distributions for 50 graphs.
Results
DiGNNExplainer achieves the highest similarity for node degree distribution across all datasets except BA-Shapes, and the highest similarity for node type distribution (NTD) and clustering coefficient (Clus.) for the real-world datasets (Table 1).
Interpretation
DiGNNExplainer generates the most realistic graphs, preserving the distributions of graphs for all datasets and node types for heterogeneous datasets. The homogeneous datasets have a single node type, hence the comparison always yields a score of 0.0.
4.3. Realism of Node Features
In this section, we identify the approach for generating the most realistic node features using cosine similarity (Node feat. (cos) in Table 2). We compare DiTabDDPM and TabDDPM in terms of node feature realism for various node feature sizes. The notations used for our feature selection experiments in the Table 2 are defined as: (i) DiTabDDPM—discrete (bag-of-words) node features generated using DiTabDDPM and continuous node features generated using TabDDPM (ii) TabDDPM—all node features generated using TabDDPM (iii) DiTabDDPM (8 discr. feat.) indicates that the node to be classified has a feature size of 8.
Implementation
(1) For discrete node features, we use one-hot encoding of word frequencies, i.e., in the range of 0 to 5 for paper nodes. (2) For VAE, we learn the graphs along with node features, which include all features, i.e., 50 for DBLP and 3066 for IMDB. In VAE, we average the cosine similarity of the node features of the node to be classified for 50 graphs.
Results
Cosine similarity (Node feat.) results are higher for DiTabDDPM compared to TabDDPM, with the highest score of 0.397 for DBLP.
Interpretation
DiTabDDPM, which learns a discrete space for discrete features, generates more realistic node features compared to the original TabDDPM, which learns a continuous space for discrete features.
4.4. Faithfulness of Explanations
@@
In Table 3, we compare the faithfulness of our approach against existing model-level explanation approaches that do not utilize actual node features. For this comparison, we use the original homogeneous datasets from the original baseline explainer implementations. Across all settings, we keep the generated graphs and ground-truth motifs fixed and vary only the node feature sizes.
In Table 2, we train DiTabDDPM, TabDDPM, and VAE to generate heterogeneous explanation graphs with varying node feature sizes (as in (Dong et al., 2021)). We assess their realism by comparing the generated features to the ground truth features (Node feat. (cos)). Moreover, we evaluate the faithfulness of the resulting explanations by applying a GNN.
Implementation
@@
We use the original model architecture for each baseline. Since the baseline explainers do not support real-valued node features or heterogeneous message passing, we convert the heterogeneous datasets DBLP and IMDB into homogeneous graphs with one-hot node types—the only setting in which all baselines can be executed. In DiGNNExplainer-GCN, we train a GCN on the converted homogeneous versions of DBLP and IMDB, as well as on MUTAG, and synthetic datasets, whereas in DiGNNExplainer, we train GraphSAGE or HAN on the original heterogeneous versions of DBLP and IMDB, and use a GCN for the other datasets. To ensure a fair comparison, following DiGNNExplainer, each baseline utilizes a different model for each graph size. We use early stopping for all trained GNNs to prevent overfitting.
Results
In Table 3, DiGNNExplainer, which utilizes real node features, achieves superior predictive and ground-truth faithfulness compared to all baseline explainers (on DBLP, IMDB, and MUTAG). DiGNNExplainer also achieves the best ground-truth faithfulness for the synthetic datasets. On DBLP and IMDB, DiGNNExplainer-GCN performs worse than DiGNNExplainer but still outperforms all baseline explainers in both metrics.
In Table 2, for DBLP, DiTabDDPM with 50 features achieves the best explanation graphs with a maximum ground-truth faithfulness of 0.882. For IMDB, we obtain similarly high maximum scores of 0.91 for both DiTabDDPM (2 discr. feat.) and TabDDPM (2 discr. feat.). The key advantage of using ground-truth faithfulness for evaluation is robustness to spurious, out-of-distribution explanations: high GNN scores can arise from out-of-distribution graphs that would never occur in practice, whereas high ground-truth faithfulness ensures that the explanatory motifs align with the true building blocks of the data and therefore better reflect real-world behavior.
Interpretation
@@ Results
(1) The explanations generated by DiGNNExplainer for our real-world datasets (Table 3) are faithful, as they achieve both high predictive faithfulness and ground-truth faithfulness. (2) Our baseline explainers yield a lower ground-truth faithfulness due to their inability to model heterogeneous structures and lack of node features. (3) The low predictive faithfulness of DiGNNExplainer on the synthetic datasets arises because the 3-layer GCN is heavily biased toward predicting the base node class. As a result, only one out of the four classes (base, house, cycle, grid) is predicted correctly in many cases, which leads to a low average predictive faithfulness across classes. (4) The low ground-truth faithfulness of VAE (Table 2) suggests VAE-based graph generation is unable to preserve the underlying heterogeneous structures. (5) The high predictive faithfulness of VAE for IMDB can be explained by the fact that the GNN is trained with all features of the dataset, where all node types except movie have continuous features, and the continuous features generated using VAE are more realistic. (6) Realistic discrete features generated by learning a discrete space in DiTabDDPM result in highly faithful explanations, but we also observe that the node features for faithful explanations are not always the most realistic, i.e., do not have the highest cosine similarity values. (7) On DBLP (Table 2), using a larger feature size produces both high predictive faithfulness and ground-truth faithfulness scores, indicating that the GNN model can be trusted. Overall, observations (5)–(7) indicate that the node feature generation technique substantially influences the quality of explanation graphs.
4.5. Experiment on Feature Selection